# Time-dependent Regularized 13-Moment Equations with Onsager Boundary Conditions in the Linear Regime

### v1

@@ 2.2. Moment equations and boundary conditions
It will become clear by its derivation that this set of boundary conditions satisfies the general form of $L^{2}$-stable boundary conditions (6). The external contribution $\boldsymbol{g}_{\mathrm{ext}}$ is given by the terms with wall velocity $v_{i}^{W}$ and wall temperature $\theta^{W}$.
@@ 3.1. Choice of variables
In (32), the constants $c_{1}^{1,1}$ and $c_{2}^{0,0}$ can be chosen as arbitrary nonzero numbers, and here we choose them such that when the collision model reduces Maxwell molecules, $\bar{q}_{i}$ and $\bar{\sigma}_{ij}$ reduce to $q_{i}$ and $\sigma_{ij}$. More precisely, we choose negative $c_{1}^{1,1}$ and positive $c_{2}^{0,0}$ such that
@@ 3.2. Second-order variables
The coefficients $c_{k}^{\ell,n}$ can be represented as functions of $\beta_{k}^{(\ell),n}$, and the details can be found in the appendix B. Here we only emphasize that the choice has been made such that $\langle\phi_{i}^{2},\phi_{i}^{3}\rangle=\langle\phi_{ij}^{1},\phi_{ij}^{2}\rangle=0$, and in the case of Maxwell molecules, $\phi_{i}^{2,3}=\psi_{i}^{2,3}$ and $\phi_{ij}^{1,2}=\psi_{ij}^{1,2}$. The result (37) indicates that the minimum number of second-order variables is 24 ($=1+3+3+5+5+7$). These variables will be denoted by
@@ 3.3. Derivation of R13 equations
The coefficients $A_{ij}$ are series of $c_{k}^{l,n}$, with their expressions provided in the appendix C.
It is now clear that the second-order variables defined in (39) can be solved according to the equations (47)–(52). Plugging the results into (43)–(46) yields our final equations (7)–(11).
@@ 3.4.3. Third Maxwellian iteration
(59) $\begin{split}&\mathcal{L}^{(11)}f_{3}^{(1)}+\mathcal{L}^{(12)}f_{3}^{(2)}+\mathcal{L}^{(1\mathrm{r})}f_{3}^{(\mathrm{r})}\\ ={}&\mathrm{Kn}\mathcal{A}^{(10)}f^{(0)}+\mathrm{Kn}^{2}\left(\mathcal{A}^{(11)}S_{1}^{(10)}-S_{1}^{(10)}\mathcal{A}^{(00)}\right)f^{(0)}\\ &{}+\mathrm{Kn}^{3}\left(\mathcal{A}^{(11)}\mathcal{S}_{2}^{(10)}+\mathcal{A}^{(12)}\mathcal{S}_{2}^{(20)}-\mathcal{S}_{2}^{(20)}\mathcal{A}^{(00)}-\mathcal{S}_{1}^{(10)}\mathcal{A}^{(01)}\mathcal{S}_{1}^{(10)}\right)f^{(0)}.\end{split}$
@@ 3.4.4. Super-Burnett equations
$\mathcal{A}^{(02)}=\mathcal{A}^{(01)}(\mathcal{L}^{(11)})^{-1}\mathcal{L}^{(12)},\qquad\mathcal{A}^{(01}=\mathcal{A}^{(01)}(\mathcal{L}^{(11)})^{-1}\mathcal{L}^{(1\mathrm{r})}.$
@@
$\frac{\partial f^{(0)}}{\partial t}+\mathcal{A}^{(00)}f^{(0)}+\mathcal{A}^{(01)}(\mathcal{L}^{(11)})^{-1}\left(\mathcal{L}^{(11)}f_{3}^{(1)}+\mathcal{L}^{(11)}f_{3}^{(2)}+\mathcal{L}^{(11)}f_{3}^{(3)}\right)=0.$
@@ 4.2. One-dimensional problems and spurious boundary layers
$\xi=\mathrm{Kn}\frac{\partial\theta}{\partial x},\qquad\eta=\mathrm{Kn}\frac{\partial\bar{\sigma}}{\partial x},$
@@
$\mathrm{Kn}\begin{pmatrix}0&0&1&0\\ 0&0&0&1\\ -\left(k_{6}+\frac{4}{5}k_{7}\right)\frac{3k_{1}}{k_{0}}&\left(k_{6}+\frac{4}{5}k_{7}\right)\frac{3k_{2}}{k_{0}}&0&0\\ 2k_{2}+\frac{4k_{8}k_{1}}{5k_{0}}&-\left(\frac{6}{5}k_{9}+\frac{2}{3}k_{10}+\frac{4k_{8}k_{2}}{5k_{0}}\right)&0&0\\ \end{pmatrix}\frac{\partial}{\partial x}\begin{pmatrix}\xi\\ \eta\\ \theta\\ \bar{\sigma}\end{pmatrix}$
$+\begin{pmatrix}-1&0&0&0\\ 0&-1&0&0\\ 0&0&\frac{5}{2}k_{0}+\frac{l_{1}k_{1}}{k_{0}}&k_{8}-\frac{l_{1}k_{2}}{k_{0}}\\ 0&0&0&l_{2}\\ \end{pmatrix}\begin{pmatrix}\xi\\ \eta\\ \theta\\ \bar{\sigma}\end{pmatrix}=\begin{pmatrix}0\\ 0\\ C_{\bar{q}}^{1}x+C_{\bar{q}}^{2}\\ 0\end{pmatrix}.$
@@
(83) $\mathrm{Kn}\frac{\partial}{\partial x}\begin{pmatrix}\xi\\ \eta\\ \theta\\ \bar{\sigma}\end{pmatrix}+\begin{pmatrix}0&0&b_{11}&b_{12}\\ 0&0&b_{21}&b_{22}\\ -1&0&0&0\\ 0&-1&0&0\end{pmatrix}\begin{pmatrix}\xi\\ \eta\\ \theta\\ \bar{\sigma}\end{pmatrix}=\begin{pmatrix}c_{1}(C_{\bar{q}}^{1}x+C_{\bar{q}}^{2})\\ c_{2}(C_{\bar{q}}^{1}x+C_{\bar{q}}^{2})\\ 0\\ 0\end{pmatrix}.$
@@
For Maxwell molecules, the above procedure of solving one-dimensional equations can be further simplified from the third step (80). By noticing $k_{1}=k_{2}=0$, one can get $\frac{\partial\bar{q}}{\partial x}=0$ and therefore (80) becomes $\bar{q}(x)=C_{\bar{q}}$. The equation (82) becomes a second-order ODE for $\bar{\sigma}$, so that the general solution of $\bar{\sigma}(x)$ contains only one boundary layer. Furthermore, (81) is in absence of second-order derivatives, and thus
@@
The thinness of the second boundary layer also indicates the largeness of the corresponding eigenvalue $\lambda_{2}$. To illustrate the magnitude of eigenvalues, we compute them according to (84) for various inverse-power-law models and summarize the results in table 1. For Maxwell molecules ($\eta=5$), only $\lambda_{1}$ is provided. The table clearly shows that all models share one similar eigenvalue $\lambda_{1}$ and each of the non-Maxwell cases has an additional larger eigenvalue $\lambda_{2}$. This eigenvalue $\lambda_{2}$ approaches infinity when $\eta$ tends to $5$.

### v2

@@ 2.2. Moment equations and boundary conditions
It will become clear by its derivation that this set of boundary conditions satisfies the general form of $L^{2}$-stable boundary conditions (6), which implies the second law of thermodynamics. The external contribution $\boldsymbol{g}_{\mathrm{ext}}$ is given by the terms with wall velocity $v_{i}^{W}$ and wall temperature $\theta^{W}$.
2.3. Second law of thermodynamics
The derivation of the new R13 equations (7)—(11), to be detailed in the next section, maintains the structure required by (2). Therefore, according to (3). The second law of thermodynamics is automatically satisfied. In fact, the entropy inequality can also be observed from (7)—(11) directly by straightforward calculations. Assume that the spatial domain is $\mathbb{R}^{3}$ and all quantities decay to zero when $\boldsymbol{x}$ tends infinity. Then, by calculating
[…]
one obtains
(22) $\begin{split}&\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}^{3}}\frac{1}{2}\left(\rho^{2}+\frac{3}{2}\theta^{2}+|\boldsymbol{v}|^{2}+\frac{2}{5}|\boldsymbol{\bar{q}}|^{2}+\frac{1}{2}\|\boldsymbol{\bar{\sigma}}\|^{2}\right)\mathrm{d}\boldsymbol{x}\\ =&-\mathrm{Kn}\int_{\mathbb{R}^{3}}\left(\frac{3}{2}k_{1}|\nabla\theta|^{2}+k_{3}\|(\nabla\boldsymbol{v})_{\mathrm{stf}}\|^{2}+\frac{4}{5}k_{6}\|\nabla\boldsymbol{\bar{q}}\|^{2}+\frac{24}{25}k_{7}\|(\nabla\boldsymbol{\bar{q}})_{\mathrm{stf}}\|^{2}\right)\mathrm{d}\boldsymbol{x}\\ &-\mathrm{Kn}\int_{\mathbb{R}^{3}}\left(k_{9}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|(\nabla\boldsymbol{\bar{\sigma}})_{\mathrm{stf}}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}^{2}+\frac{1}{2}k_{10}|\nabla\cdot\boldsymbol{\bar{\sigma}}|^{2}\right)\mathrm{d}\boldsymbol{x}-\frac{1}{\mathrm{Kn}}\int_{\mathbb{R}^{3}}\left(\frac{4}{15}l_{1}|\boldsymbol{\bar{q}}|^{2}+\frac{1}{2}l_{2}\|\boldsymbol{\bar{\sigma}}\|^{2}\right)\mathrm{d}\boldsymbol{x}\\ &+3\mathrm{Kn}k_{2}\int_{\mathbb{R}^{3}}\nabla\theta\cdot\left(\nabla\cdot\boldsymbol{\bar{\sigma}}\right)\mathrm{d}\boldsymbol{x}-2\mathrm{Kn}k_{4}\int_{\mathbb{R}^{3}}\left(\nabla\boldsymbol{v}\right)_{\mathrm{stf}}:\left(\nabla\boldsymbol{\bar{q}}\right)_{\mathrm{stf}}\mathrm{d}\boldsymbol{x},\end{split}$
[…]
(23) $\frac{(3k_{2})^{2}}{4}\leqslant\frac{3}{2}k_{1}\cdot\frac{1}{2}k_{10},\qquad k_{4}^{2}\leqslant k_{3}\cdot\frac{24}{25}k_{7},$
which yields
$\left|3k_{2}\int_{\mathbb{R}^{3}}\nabla\theta\cdot(\nabla\cdot\boldsymbol{\bar{\sigma}})\,\mathrm{d}\boldsymbol{x}\right|\leqslant\int_{\mathbb{R}^{3}}\left(\frac{3}{2}k_{1}|\nabla\theta|^{2}+\frac{1}{2}k_{10}|\nabla\cdot\boldsymbol{\bar{\sigma}}|^{2}\right)\mathrm{d}\boldsymbol{x},$
$\left|2k_{4}\int_{\mathbb{R}^{3}}(\nabla\boldsymbol{v})_{\mathrm{stf}}:(\nabla\boldsymbol{\bar{q}})_{\mathrm{stf}}\,\mathrm{d}\boldsymbol{x}\right|\leqslant\int_{\mathbb{R}^{3}}\left(k_{3}\|(\nabla\boldsymbol{v})_{\mathrm{stf}}\|^{2}+\frac{24}{25}k_{7}\|(\nabla\boldsymbol{\bar{q}})_{\mathrm{stf}}\|^{2}\right)\mathrm{d}\boldsymbol{x}.$
Thus,
(24) $\begin{split}&\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}^{3}}\frac{1}{2}\left(\rho^{2}+\frac{3}{2}\theta^{2}+|\boldsymbol{v}|^{2}+\frac{2}{5}|\boldsymbol{\bar{q}}|^{2}+\frac{1}{2}\|\boldsymbol{\bar{\sigma}}\|^{2}\right)\mathrm{d}\boldsymbol{x}\\ \leqslant{}&{-\mathrm{Kn}}\int_{\mathbb{R}^{3}}\left(\frac{4}{5}k_{6}\|\nabla\boldsymbol{\bar{q}}\|^{2}+k_{9}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|(\nabla\boldsymbol{\bar{\sigma}})_{\mathrm{stf}}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}^{2}\right)\mathrm{d}\boldsymbol{x}-\frac{1}{\mathrm{Kn}}\int_{\mathbb{R}^{3}}\left(\frac{4}{15}l_{1}|\boldsymbol{\bar{q}}|^{2}+\frac{1}{2}l_{2}\|\boldsymbol{\bar{\sigma}}\|^{2}\right)\mathrm{d}\boldsymbol{x},\end{split}$
Since the coefficients $k_{6}$, $k_{9}$, $l_{1}$ and $l_{2}$ are all nonnegative, the right-hand side of the equation above is nonpositive. The entropy density can thus be defined by
$H=H_{0}-\frac{1}{2}\left(\rho^{2}+\frac{3}{2}\theta^{2}+|\boldsymbol{v}|^{2}+\frac{2}{5}|\boldsymbol{\bar{q}}|^{2}+\frac{1}{2}\|\boldsymbol{\bar{\sigma}}\|^{2}\right),$
where $H_{0}$ is an arbitrary constant. In the case of Maxwell molecules, this results echos the conclusion in [35].
For bounded domains, with boundary conditions (13)—(21), we can also show that the entropy may increase only due to the incoming fluxes from the boundary. The derivation is similar to the unbounded case, while boundary terms appear when performing integration by parts, for which the boundary conditions need to be applied to make further simplifications. For any bounded domain $\Omega$, the conclusion has the following form:
(25) $\begin{split}-\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}H\,\mathrm{d}\boldsymbol{x}\leqslant{}&\int_{\partial\Omega}\theta^{W}(C_{1}\theta+C_{2}\bar{q}_{n}+C_{3}\bar{\sigma}_{nn})\,\mathrm{d}s\\ +\int_{\partial\Omega}[v_{\tau_{1}}^{W}&(C_{4}v_{\tau_{1}}+C_{5}\bar{q}_{\tau_{1}}+C_{6}\bar{\sigma}_{n\tau_{1}})+v_{\tau_{2}}^{W}(C_{4}v_{\tau_{2}}+C_{5}\bar{q}_{\tau_{2}}+C_{6}\bar{\sigma}_{n\tau_{2}})]\,\mathrm{d}s,\end{split}$
where the constants $C_{1}$ to $C_{6}$ depend only on the coefficients in the equations and the boundary conditions. In (25), the right-hand side only contains terms related to $\boldsymbol{v}^{W}$ and $\theta^{W}$, meaning that the entropy can increase only due to the velocities and temperatures of the walls, which does not violate the second law of thermodynamics.
@@ 3.1. Choice of variables
In (38), the constants $c_{1}^{1,1}$ and $c_{2}^{0,0}$ can be chosen as arbitrary nonzero numbers, and here we choose them such that when the collision model reduces Maxwell molecules ($\eta=5$), $\bar{q}_{i}$ and $\bar{\sigma}_{ij}$ reduce to $q_{i}$ and $\sigma_{ij}$. More precisely, we choose negative $c_{1}^{1,1}$ and positive $c_{2}^{0,0}$ such that
@@ 3.2. Second-order variables
The coefficients $c_{k}^{\ell,n}$ can be represented as functions of $\beta_{k}^{(\ell),n}$, and the details can be found in the appendix D. Here we only emphasize that the choice has been made such that $\langle\phi_{i}^{2},\phi_{i}^{3}\rangle=\langle\phi_{ij}^{1},\phi_{ij}^{2}\rangle=0$, and in the case of Maxwell molecules, $\phi_{i}^{2,3}=\psi_{i}^{2,3}$ and $\phi_{ij}^{1,2}=\psi_{ij}^{1,2}$. The result (43) indicates that the minimum number of second-order variables is 24 ($=1+3+3+5+5+7$). These variables will be denoted by
@@ 3.3. Derivation of R13 equations
The coefficients $A_{ij}$ are series of $c_{k}^{l,n}$, with their expressions provided in the appendix E.
It is now clear that the second-order variables defined in (45) can be solved according to the equations (53)–(58). Plugging the results into (49)–(52) yields our final equations (7)–(11). Here we would like to write down the explicit expressions for a few coefficients in (7)–(11):
$k_{1}=-5\begin{pmatrix}c_{1}^{2,1}&c_{1}^{3,1}\end{pmatrix}\begin{pmatrix}\mathscr{L}_{1}^{(22)}&\mathscr{L}_{1}^{(23)}\\ \mathscr{L}_{1}^{(32)}&\mathscr{L}_{1}^{(33)}\end{pmatrix}^{-1}\begin{pmatrix}c_{1}^{2,1}\\ c_{1}^{3,1}\end{pmatrix},$
$k_{2}=-\frac{\sqrt{2}}{3}\begin{pmatrix}c_{1}^{2,1}&c_{1}^{3,1}\end{pmatrix}\begin{pmatrix}\mathscr{L}_{1}^{(22)}&\mathscr{L}_{1}^{(23)}\\ \mathscr{L}_{1}^{(32)}&\mathscr{L}_{1}^{(33)}\end{pmatrix}^{-1}\begin{pmatrix}A_{57}\\ A_{58}\end{pmatrix},$
$k_{3}=-15\begin{pmatrix}c_{2}^{1,0}&c_{2}^{2,0}\end{pmatrix}\begin{pmatrix}\mathscr{L}_{2}^{(11)}&\mathscr{L}_{2}^{(12)}\\ \mathscr{L}_{2}^{(21)}&\mathscr{L}_{2}^{(22)}\end{pmatrix}^{-1}\begin{pmatrix}c_{2}^{1,0}\\ c_{2}^{2,0}\end{pmatrix},$
$k_{4}=\sqrt{2}\begin{pmatrix}A_{49}&A_{4,10}\end{pmatrix}\begin{pmatrix}\mathscr{L}_{2}^{(11)}&\mathscr{L}_{2}^{(12)}\\ \mathscr{L}_{2}^{(21)}&\mathscr{L}_{2}^{(22)}\end{pmatrix}^{-1}\begin{pmatrix}c_{2}^{1,0}\\ c_{2}^{2,0}\end{pmatrix},$
$k_{7}=-\frac{5}{36}\begin{pmatrix}A_{49}&A_{4,10}\end{pmatrix}\begin{pmatrix}\mathscr{L}_{2}^{(11)}&\mathscr{L}_{2}^{(12)}\\ \mathscr{L}_{2}^{(21)}&\mathscr{L}_{2}^{(22)}\end{pmatrix}^{-1}\begin{pmatrix}A_{49}\\ A_{4,10}\end{pmatrix},$
$k_{10}=-\frac{2}{15}\begin{pmatrix}A_{57}&A_{58}\end{pmatrix}\begin{pmatrix}\mathscr{L}_{1}^{(22)}&\mathscr{L}_{1}^{(23)}\\ \mathscr{L}_{1}^{(32)}&\mathscr{L}_{1}^{(33)}\end{pmatrix}^{-1}\begin{pmatrix}A_{57}\\ A_{58}\end{pmatrix}.$
Since the matrices
$\begin{pmatrix}\mathscr{L}_{1}^{(22)}&\mathscr{L}_{1}^{(23)}\\ \mathscr{L}_{1}^{(32)}&\mathscr{L}_{1}^{(33)}\end{pmatrix}\quad\text{and}\quad\begin{pmatrix}\mathscr{L}_{2}^{(11)}&\mathscr{L}_{2}^{(12)}\\ \mathscr{L}_{2}^{(21)}&\mathscr{L}_{2}^{(22)}\end{pmatrix}$
are symmetric and negative definite, we can obtain (23) by the Cauchy-Schwarz inequality, which completes the proof of the H-theorem (24).
@@ 3.4.3. Third Maxwellian iteration
(65) $\begin{split}&\mathcal{L}^{(11)}f_{3}^{(1)}+\mathcal{L}^{(12)}f_{3}^{(2)}+\mathcal{L}^{(1\mathrm{r})}f_{3}^{(\mathrm{r})}\\ ={}&\mathrm{Kn}\mathcal{A}^{(10)}f^{(0)}+\mathrm{Kn}^{2}\left(\mathcal{A}^{(11)}\mathcal{S}_{1}^{(10)}-\mathcal{S}_{1}^{(10)}\mathcal{A}^{(00)}\right)f^{(0)}\\ &{}+\mathrm{Kn}^{3}\left(\mathcal{A}^{(11)}\mathcal{S}_{2}^{(10)}+\mathcal{A}^{(12)}\mathcal{S}_{2}^{(20)}-\mathcal{S}_{2}^{(20)}\mathcal{A}^{(00)}-\mathcal{S}_{1}^{(10)}\mathcal{A}^{(01)}\mathcal{S}_{1}^{(10)}\right)f^{(0)}.\end{split}$
@@ 3.4.4. Super-Burnett equations
$\mathcal{A}^{(02)}=\mathcal{A}^{(01)}(\mathcal{L}^{(11)})^{-1}\mathcal{L}^{(12)},\qquad\mathcal{A}^{(0\mathrm{r})}=\mathcal{A}^{(01)}(\mathcal{L}^{(11)})^{-1}\mathcal{L}^{(1\mathrm{r})}.$
@@
$\frac{\partial f^{(0)}}{\partial t}+\mathcal{A}^{(00)}f^{(0)}+\mathcal{A}^{(01)}(\mathcal{L}^{(11)})^{-1}\left(\mathcal{L}^{(11)}f_{3}^{(1)}+\mathcal{L}^{(12)}f_{3}^{(2)}+\mathcal{L}^{(1\mathrm{r})}f_{3}^{(\mathrm{r})}\right)=0.$
@@ 4. Onsager boundary conditions
Remark 2.
Note that existing Onsager boundary conditions such as the ones derived in [29] and [10] cannot be directly applied to our systems. The primary reason is that our system generally requires more boundary conditions than previous R13 equations, due to a larger number of moments considered in our derivation. For instance, the boundary conditions for R13 equations with Maxwell molecules consist of 10 equations at each boundary point [29]. However, in our system, the equations of $\theta$ and $\boldsymbol{v}$ become parabolic, requiring 3 more boundary conditions to determine the solution. This motivates us to derive the boundary conditions from scratch and guarantee reliable matching between the equations and boundary conditions.
@@ 4.2. One-dimensional problems and spurious boundary layers
$\xi=\mathrm{Kn}\frac{\partial\theta}{\partial x},\qquad\xi^{\prime}=\mathrm{Kn}\frac{\partial\bar{\sigma}}{\partial x},$
@@
$\mathrm{Kn}\begin{pmatrix}0&0&1&0\\ 0&0&0&1\\ -\left(k_{6}+\frac{4}{5}k_{7}\right)\frac{3k_{1}}{k_{0}}&\left(k_{6}+\frac{4}{5}k_{7}\right)\frac{3k_{2}}{k_{0}}&0&0\\ 2k_{2}+\frac{4k_{8}k_{1}}{5k_{0}}&-\left(\frac{6}{5}k_{9}+\frac{2}{3}k_{10}+\frac{4k_{8}k_{2}}{5k_{0}}\right)&0&0\\ \end{pmatrix}\frac{\partial}{\partial x}\begin{pmatrix}\xi\\ \xi^{\prime}\\ \theta\\ \bar{\sigma}\end{pmatrix}$
$+\begin{pmatrix}-1&0&0&0\\ 0&-1&0&0\\ 0&0&\frac{5}{2}k_{0}+\frac{l_{1}k_{1}}{k_{0}}&k_{8}-\frac{l_{1}k_{2}}{k_{0}}\\ 0&0&0&l_{2}\\ \end{pmatrix}\begin{pmatrix}\xi\\ \xi^{\prime}\\ \theta\\ \bar{\sigma}\end{pmatrix}=\begin{pmatrix}0\\ 0\\ C_{\bar{q}}^{1}x+C_{\bar{q}}^{2}\\ 0\end{pmatrix}.$
@@
(89) $\mathrm{Kn}\frac{\partial}{\partial x}\begin{pmatrix}\xi\\ \xi^{\prime}\\ \theta\\ \bar{\sigma}\end{pmatrix}+\begin{pmatrix}0&0&b_{11}&b_{12}\\ 0&0&b_{21}&b_{22}\\ -1&0&0&0\\ 0&-1&0&0\end{pmatrix}\begin{pmatrix}\xi\\ \xi^{\prime}\\ \theta\\ \bar{\sigma}\end{pmatrix}=\begin{pmatrix}c_{1}(C_{\bar{q}}^{1}x+C_{\bar{q}}^{2})\\ c_{2}(C_{\bar{q}}^{1}x+C_{\bar{q}}^{2})\\ 0\\ 0\end{pmatrix}.$
@@
For Maxwell molecules ($\eta=5$), the above procedure of solving one-dimensional equations can be further simplified from the third step (86). By noticing $k_{1}=k_{2}=0$, one can get $\frac{\partial\bar{q}}{\partial x}=0$ and therefore (86) becomes $\bar{q}(x)=C_{\bar{q}}$. The equation (88) becomes a second-order ODE for $\bar{\sigma}$, so that the general solution of $\bar{\sigma}(x)$ contains only one boundary layer. Furthermore, (87) is in absence of second-order derivatives, and thus
@@
The thinness of the second boundary layer also indicates the largeness of the corresponding eigenvalue $\lambda_{2}$. To illustrate the magnitude of eigenvalues, we compute them according to (90) for various inverse-power-law models and summarize the results in table 1. For Maxwell molecules, only $\lambda_{1}$ is provided. The table clearly shows that all models share one similar eigenvalue $\lambda_{1}$ and each of the non-Maxwell cases has an additional larger eigenvalue $\lambda_{2}$. This eigenvalue $\lambda_{2}$ approaches infinity when $\eta$ tends to $5$.